# Learning Zero-Shot Material States Segmentation, by Implanting Natural Image Patterns in Synthetic Data

### v5

@@ Abstract
Figure 1: Top: MatSeg benchmark images and MatSeg trained net predictions. Predictions are shown as a material similarity map relative to a selected point in the image (marked green). Center: Synthetic 3D scenes and their annotations. Different colors in the annotation stand for different materials, black is the background. Bottom: Textures and SVBRDF/PBR materials automatically extracted from natural images. More samples can be seen at figs. 10, 11 and 7.
@@ 1 Introduction
Materials and their states form a vast array of patterns and textures that define the physical and visual world. Minerals in rocks, sediment in soil, dust on surfaces, infection on leaves, stains on fruits, and foam in liquids are some of these almost infinite numbers of states and patterns. Segmenting these states in images is fundamental to the comprehension of the world and is essential for a wide range of tasks, from cooking and cleaning to construction and laboratory work. Currently, there is no comprehensive dataset or benchmark that addresses this general task, and even the top foundation models perform poorly on this problem (Section 6). This work presents the first general benchmark and datasets focused on this task, encompassing a wide range of material states and domains without being limited to specific materials or settings[7, 59]. This includes dealing with gradual transitions and partial similarity between states, as well as scattered shapes and fuzzy boundaries (Figure 1 top). For the benchmark, the goal is to achieve this using real-world images taken in the wild. The goal of the training set is to provide unlimited synthetic data that capture the complexity and variability of the real world while not being limited to a set of procedural rules and assets. Manual annotation of material states is highly complex from a human perspective, especially for cases with soft boundaries and gradual transitions or sparse and scattered patterns. In specific settings, it is sometimes possible to find some method or additive to simplify the annotations. However, the unstructured and wide range of domains in the benchmark makes this unattainable in this case. Synthetic data can simulate soft and scattered states with perfect annotation. However, it fails to capture the vast complexity of material forms in the real world, and while it is possible to accurately represent specific domains using simulation or procedural rules[45]. Achieving this for the full range of material states in the world encompasses the entire material science and is unattainable.
@@ 1.1 Procedural imitation: Extracting and implanting real-world patterns in synthetic data
This work offers a method to bridge this gap between synthetic and real data by automatically extracting patterns and textures from real-world images and using these to generate and map materials into synthetic scenes (Figure 2). This approach combines the precision and scale of synthetic data with the vast variability of patterns in the real world. The hypothesis is that simple image properties like value, hue, saturation, or one of the Red, Green, or Blue channels (Figure 2) are often correlated with material regions or properties within a material texture[15, 63, 27]. For example, the dry part of a leaf, stains on fabrics, or pollution in water will often be darker or brighter than their surroundings. In these cases, the image brightness map will capture the material region. Similarly, within a single material texture, regions with higher reflectivity, transparency, or roughness will often be darker or brighter, allowing a single image channel to capture the property distribution. Assuming this is true for some fraction of cases and that important material states appear in a wide range of settings, extracting these maps from a broad range of images will capture these patterns. Once captured, these patterns can be used in various synthetic scenes where the original correlation is not applied and therefore will be learned by the net in general form (Figure 2). Thus, patterns are automatically extracted in simple cases with a strong correlation between material and a simple image property but are then rendered and learned in other scenes where this correlation does not occur. Since this is an unsupervised process, it will capture a significant amount of noise and unrelated patterns. However, neural nets are very effective in learning from noisy data, as demonstrated by methods such as CLIP and GPT[52, 44, 9]. Conversely, neural nets tend to perform poorly when trained on clean data with a narrow domain and then are expected to generalize to broader tasks[44, 29]. The fact that the net trained on this infused data outperforms top foundation models (SAM, Materialistic[34, 53]) on a wide range of benchmarks strongly supports this hypothesis. We share the dataset and an extensive repository of extracted textures and SVBRDF/PBR materials collected by this method to enable the generation of future datasets (appendix A).
@@ 1.2 Zeros-Shot Material State Segmentation Benchmark
We present the first general benchmark for zero-shot material state segmentation. Benchmarks for specific material states like corrosion or dust have been published[48, 18]; however, no general zero-shot benchmark for material states that encompasses a wide range of fields without being limited to a specific set of states has been published so far. The benchmark contains 1,220 real-world images with a wide range of materials and domains, such as food states (cooked / burned, etc.), plants (infected / dry), rocks / soil (minerals / sediment), construction / metals (rusted/worn), liquids (foam/precipitate), and many other states in a class-agnostic manner (figs. 1 and 3). The goal is to evaluate the segmentation of material states without being limited to a specific class, domain, or setting. This includes materials with complex or scattered shapes, fuzzy boundaries, or gradual transitions. This makes manual polygon-based annotation unfeasible. To solve this, we use point-based and similarity-based annotations. Hence, for each image, we sample multiple points that represent the distribution of each material state (fig. 3). Points with the exact same material state are grouped under the same label. We also define groups (materials) that have partial similarity. For example, points in group A are more similar to points in group B than to points in group C if materials A and B are similar but not identical. This captures the complexity of gradual transitions while also dealing with scattered shapes without needing to manually annotate the full image, which in many cases is unfeasible. The top image segmentation foundation models (SAM, Materialistic[34, 53]) perform poorly on this benchmark. This shows that standard data collection approaches have a major gap for these types of tasks.
@@ 3 Data generation: patterns extraction and infusion
The data infusion process described in figs. 1 and 1.1 serves two main purposes. The first is to map existing materials into synthetic scenes (section 3.1). The second application involves extracting texture maps and PBR/SVBRDF materials from images (section 3.2). Both applications rely on the assumptions outlined in section 1.1 and are detailed below.
@@ 3.1 Unsupervised extraction of maps from images
The extraction of maps from images was done by randomly picking one property of the image; this can be the saturation, hue, brightness, or one of the RGB channels (fig. 2b). Each of these features transforms the image into a 2D topological map. The selected map passes through a fuzzy threshold (color ramp, fig. 2c). This involves picking two random threshold values. Everything above the higher threshold turns 1, everything below the lower threshold turns 0, and everything in between is a linear interpolation between 0 and 1. This allows the map to have distinct segments while maintaining soft and gradual transitions. Repeating this process several times can create maps with several distinct regions. This map can be used to map materials onto the surface of objects and scenes (fig. 4).
Creating 2D Scenes: The simple way to use the extracted map is to define every map region as a different material in an image and map different textures into different regions (Figure 2.center.d). The resulting image is then overlayed on top of a random background image (Figure 2.center.e). Shadow-like effects can be achieved by using another extracted soft map to darken some regions according to the map value. In regions where the value of the map is between 0 and 1, several textures are mixed according to the values of the map (fig. 4).
Figure 4: Examples of 2D and 3D synthetic scenes (top) and their annotations (bottom). In the annotation, each material is marked as a different color. A mixture of materials marked as a mixture of their colors. The background is marked black. More samples can be seen at fig. 11.
Creating 3D scenes: A more advanced approach is to use the generated maps to set PBR materials on object surfaces in 3D scenes using CGI software such as Blender[8]. This is done by creating 3D scenes and wrapping the extracted maps around objects’ surfaces in a process called UV mapping. Random PBR materials are then placed in each segment of the map (fig. 2.left.d,e). Background, random objects, ground, and illumination are added to generate full 3D scenes with physics-based rendered materials (PBR), natural illumination, shadows, and curved surfaces (fig. 1.center).
@@ 3.2 Extracting Textures and Materials
Materials in synthetic datasets are usually represented as SVBRDF/PBR, in which the spatial distribution of each material property (reflectivity, roughness, transparency, etc.) is given as a 2D map[6, 43]. Extracting textures and PBR materials from images involves first identifying regions with a uniform texture (which implies a uniform material, fig. 2.right). This texture image is then used to extract maps corresponding to each material’s property.
@@ 3.2.1 Unsupervised texture extraction
Unsupervised texture extraction is traditionally done by finding image regions with a uniform distribution of color and gradients[57, 12, 42, 31, 62]. Finding these regions is done by splitting the image into square cells of 40x40 pixels each and identifying square regions where the distribution of the RGB values and their gradients is similar for all the cells in the region. Regions of more than 6x6 cells with similar distributions between all cells were extracted (fig. 2.right). More accurately, a region was considered to have uniform texture if the similarity of the distributions between each pair of cells in the region had a Jensen-Shannon distance of less than 0.5. This similarity was tested separately for each of the RGB channels and their gradients. Finally, regions with too-uniform values or very high or low values are ignored, as they only contain smooth, textureless regions. It’s clear that this approach will fail to capture textures on bent surfaces or non-uniform illumination and shadows. However, the assumption is that for a large enough set of images, common material textures will appear in some cases in a simple setting and will be captured (section 1.1). Once captured, they will be projected into scenes with different illuminations, and bent surfaces and will be learned for the general case. Some textures extracted using this approach are shown in figs. 1 and 8.
@@ 3.2.2 Generating material property maps from texture images
Extracting PBR material maps from RGB textures requires guessing the physical properties of the material at each point of the texture image. There is no straightforward way to achieve this using an RGB image. Once again, the assumption will be that the properties of the material (reflectivity, roughness, and transparency) are correlated, in some cases, with simple image properties like color(section 1.1), and for a common materials and a large enough set of images such correlation is bound to occur in some setting. Each material property map is generated by randomly choosing one of the six image channels (RGB or HSV), to represent this map (fig. 2 left), image augmentation, thresholding, and random scaling/shifting were also used. Normal maps were generated by taking the gradient of the height/bumpiness map(fig. 2.left). In some cases, the map was chosen to have a uniform random value. Some results are shown in figs. 1 and 7. In addition, materials were also generated by mixing maps of existing PBR materials as described in appendix F.
@@ 4 Material State Segmentation Benchmark
Image collection and annotation: 1220 images for the benchmark were manually taken from a wide variety of real-world scenes (fig. 3). Including food at different levels of cooking, liquids, solutes, foams, plants, rocks, minerals, sediments, grounds, landscapes, construction sites, metals, relics, fabrics, labs and processes like cooking, wetting, rotting, infections, corrosion, stains, and many others. The goal is to capture as much of the world as possible, both in terms of materials and environments. For materials with scatter, soft boundaries, and gradual transitions, it is often impossible to segment the entire region belonging to a material state. An alternative approach is to sample points in the image that represent the main materials’ states and their distribution. This allows annotators to focus on regions with clear annotation, but also to sample the harder and more complex regions without dealing with areas where the segmentation is unclear. The annotation procedure is as follows: 1) Pick points in each image that represent the distribution of each material state (fig. 3). 2) Group the points according to their material state such that all points belonging to the exact same material state will have the same label or group (points of the same color in fig. 3 are in the same group). 3) Assign relative similarity between groups. For example, if material B is a transition state between materials A and C, we can define the points in group B to be partially similar to both groups A and C, while groups A and C are not similar to each other. This circumvents the almost impossible task of assigning numerical similarity to material states (fig. 3).
@@ 4.1 Evaluation metric
The evaluation approach is based on assessing the relative similarity between the points in the image. Every segmentation approach, soft or hard, can be easily converted to predict the similarity between two points in the image. For hard segmentation, this similarity will be one if the two points are in the same segment and zero if not. For evaluation, we use the triplet metric which goes as follows: a) Select three points in the image and define one as an anchor. b) Ask which of the remaining two points is more similar to the anchor according to the ground truth (GT) and according to the prediction; if the predictions and GT agree, we assign this as correct. c) If the two points have identical similarity to the anchor (according to the GT), we ignore this triplet. This is done for all sets of GT triplets in the image, and the average per material is used. For this metric, a 50% score means a random guess and a 100% score means perfect segmentation. The results are given in table 1.
@@ 4.2 Additional benchmarks
While there are no other benchmarks for general zero-shot material state segmentation, there are a number of benchmarks that deal with subdomains of this task. These include segmentation of specific material states, like corrosion[48], minerals/ores[17, 16, 26, 28], dust[18], leaf disease[5], soil types and states[49, 50], microscopy[39, 40], or material phases (liquids, solids, foams, or powders)[24, 23]. Although each of these benchmarks is relatively narrow, when combined, they offer a way to cross-check the results of the main benchmark. Materialistic is a benchmark[53] that contains 50 images with different materials, not limited to specific classes. With materials segments that cover the whole object or regions with straight, smooth boundaries. It consists of distinct material types with no states or gradual transitions. Comparing the results of this benchmark to MatSeg offers a way to examine how much these elements affect segmentation accuracy. To evaluate the nets on these benchmarks, we select a random point in each material segment and use the nets to predict the similarity of each pixel in the image to the selected point. To turn this map into a hard segmentation map, we choose a threshold that maximizes the IOU between the predicted and GT masks. We use the mean IOU as the benchmark score. This is done with multiple points to obtain accurate statistics. The results are given in table 2.
@@ 6 Results
Results for the MatSeg benchmark are given in table 1 and fig. 5. The net trained on the MatSeg data achieves good accuracy and significantly outperforms both SAM and Materialistic on both hard and soft cases. It can be seen from fig. 5 that the net learned to predict highly scattered and complex patterns with soft and gradual transitions and achieves this for a wide range of domains. Hence, despite learning from synthetic data, the net manages to generalize for a wide range of real-world cases. In addition, it seems to be able to mitigate reflections and shadows fig. 5. Both Segment Anything (SAM) and Materialistic nets performed far worse on the benchmark (table 1). From fig. 5, it seems that both SAM and Materialistic search for the boundary of the material based on the boundary of the object or bulk regions and struggle with scattered shapes and soft boundaries. SAM and Materialistic also seem to be less sensitive to the material state and more focused on matching materials based on coarse-grain type (fig. 5). Both nets seem to search for bulk regions with smooth boundaries, which are likely to be more representative in their training set. SAM has failed to achieve high accuracy even after receiving 8 guiding input points (compared to 1 for other methods), suggesting that despite training on a vast amount of examples, it never encountered or learned such patterns. This seems to imply a major gap in the data collected for both SAM and Materialistic. This hypothesis is further supported by the fact that the SAM and Materialistic nets outperformed the MatSeg-trained net on the Materialistic test set (table 2), which contains distinct materials that encompass either a full object or distinct regions with smooth hard boundaries. As such, their shape distribution resembles that of assets used in the Materialistic train set and the SAM polygon-based annotations. Hence, the MatSeg data seems to better model raw material states, while SAM and Materialistic seem to perform better on more structured scenes and object-guided segments.
Figure 5: Results: Predicted material similarity maps to a given point in the image. The input point is marked green in each panel. For SAM net the results are with 8 input points, 4 positive (marked green) and 4 negative (marked red). The materials similarity map is brighter for high similarity and dark for low similarity (similarity to the input point). More results can be seen at fig. 10.
Table 1: Results on the MatSeg Benchmark (section 4). The results are in triplet loss(section 4.1). Points (pts) are the number of input points used to guide the segmentation. MatSeg 2D and 3D stand for nets trained on only 2D or 3D scenes, and Mixed stands for a net trained on a combination of 3D and 2D scenes. All stands for evaluation on all triplets. Soft stand for evaluation only on triplets with partial similarity between points(section 4).
@@
Table 2: Results on various of datasets (section 4.2). The results are in IOU for an optimal threshold (section 4.2). Benchmarks with Scattered shapes and Gradual transitions are marked (S/G). Points (pts) is the number of input points used to guide the segmentation. The top one-shot (one point) method is marked.

### v6

@@ Abstract
Figure 1: Top: MatSeg benchmark images and MatSeg trained net predictions. Predictions are shown as a material similarity map relative to a selected point in the image (marked green). Center: Synthetic 3D scenes and their annotations. Different colors in the annotation stand for different materials, black is the background. Bottom: Textures and SVBRDF/PBR materials automatically extracted from natural images. More samples can be seen at figs.˜10, 11 and 7.
@@ 1 Introduction
Materials and their states form a vast array of patterns and textures that define the physical and visual world. Minerals in rocks, sediment in soil, dust on surfaces, infection on leaves, stains on fruits, and foam in liquids are some of these almost infinite numbers of states and patterns. Segmenting these states in images is fundamental to the comprehension of the world and is essential for a wide range of tasks, from cooking and cleaning to construction and laboratory work. Currently, there is no comprehensive dataset or benchmark that addresses this general task, and even the top foundation models perform poorly on this problem (Section˜6). This work presents the first general benchmark and datasets focused on this task, encompassing a wide range of material states and domains without being limited to specific materials or settings[7, 59]. This includes dealing with gradual transitions and partial similarity between states, as well as scattered shapes and fuzzy boundaries (Figure˜1 top). For the benchmark, the goal is to achieve this using real-world images taken in the wild. The goal of the training set is to provide unlimited synthetic data that capture the complexity and variability of the real world while not being limited to a set of procedural rules and assets. Manual annotation of material states is highly complex from a human perspective, especially for cases with soft boundaries and gradual transitions or sparse and scattered patterns. In specific settings, it is sometimes possible to find some method or additive to simplify the annotations. However, the unstructured and wide range of domains in the benchmark makes this unattainable in this case. Synthetic data can simulate soft and scattered states with perfect annotation. However, it fails to capture the vast complexity of material forms in the real world, and while it is possible to accurately represent specific domains using simulation or procedural rules[45]. Achieving this for the full range of material states in the world encompasses the entire material science and is unattainable.
@@ 1.1 Procedural imitation: Extracting and implanting real-world patterns in synthetic data
This work offers a method to bridge this gap between synthetic and real data by automatically extracting patterns and textures from real-world images and using these to generate and map materials into synthetic scenes (Figure˜2). This approach combines the precision and scale of synthetic data with the vast variability of patterns in the real world. The hypothesis is that simple image properties like value, hue, saturation, or one of the Red, Green, or Blue channels (Figure˜2) are often correlated with material regions or properties within a material texture[15, 63, 27]. For example, the dry part of a leaf, stains on fabrics, or pollution in water will often be darker or brighter than their surroundings. In these cases, the image brightness map will capture the material region. Similarly, within a single material texture, regions with higher reflectivity, transparency, or roughness will often be darker or brighter, allowing a single image channel to capture the property distribution. Assuming this is true for some fraction of cases and that important material states appear in a wide range of settings, extracting these maps from a broad range of images will capture these patterns. Once captured, these patterns can be used in various synthetic scenes where the original correlation is not applied and therefore will be learned by the net in general form (Figure˜2). Thus, patterns are automatically extracted in simple cases with a strong correlation between material and a simple image property but are then rendered and learned in other scenes where this correlation does not occur. Since this is an unsupervised process, it will capture a significant amount of noise and unrelated patterns. However, neural nets are very effective in learning from noisy data, as demonstrated by methods such as CLIP and GPT[52, 44, 9]. Conversely, neural nets tend to perform poorly when trained on clean data with a narrow domain and then are expected to generalize to broader tasks[44, 29]. The fact that the net trained on this infused data outperforms top foundation models (SAM, Materialistic[34, 53]) on a wide range of benchmarks strongly supports this hypothesis. We share the dataset and an extensive repository of extracted textures and SVBRDF/PBR materials collected by this method to enable the generation of future datasets (appendix˜A).
@@ 1.2 Zeros-Shot Material State Segmentation Benchmark
We present the first general benchmark for zero-shot material state segmentation. Benchmarks for specific material states like corrosion or dust have been published[48, 18]; however, no general zero-shot benchmark for material states that encompasses a wide range of fields without being limited to a specific set of states has been published so far. The benchmark contains 1,220 real-world images with a wide range of materials and domains, such as food states (cooked / burned, etc.), plants (infected / dry), rocks / soil (minerals / sediment), construction / metals (rusted/worn), liquids (foam/precipitate), and many other states in a class-agnostic manner (figs.˜1 and 3). The goal is to evaluate the segmentation of material states without being limited to a specific class, domain, or setting. This includes materials with complex or scattered shapes, fuzzy boundaries, or gradual transitions. This makes manual polygon-based annotation unfeasible. To solve this, we use point-based and similarity-based annotations. Hence, for each image, we sample multiple points that represent the distribution of each material state (fig.˜3). Points with the exact same material state are grouped under the same label. We also define groups (materials) that have partial similarity. For example, points in group A are more similar to points in group B than to points in group C if materials A and B are similar but not identical. This captures the complexity of gradual transitions while also dealing with scattered shapes without needing to manually annotate the full image, which in many cases is unfeasible. The top image segmentation foundation models (SAM, Materialistic[34, 53]) perform poorly on this benchmark. This shows that standard data collection approaches have a major gap for these types of tasks.
@@ 3 Data generation: patterns extraction and infusion
The data infusion process described in figs.˜1 and 1.1 serves two main purposes. The first is to map existing materials into synthetic scenes (section˜3.1). The second application involves extracting texture maps and PBR/SVBRDF materials from images (section˜3.2). Both applications rely on the assumptions outlined in section˜1.1 and are detailed below.
@@ 3.1 Unsupervised extraction of maps from images
The extraction of maps from images was done by randomly picking one property of the image; this can be the saturation, hue, brightness, or one of the RGB channels (fig.˜2b). Each of these features transforms the image into a 2D topological map. The selected map passes through a fuzzy threshold (color ramp, fig.˜2c). This involves picking two random threshold values. Everything above the higher threshold turns 1, everything below the lower threshold turns 0, and everything in between is a linear interpolation between 0 and 1. This allows the map to have distinct segments while maintaining soft and gradual transitions. Repeating this process several times can create maps with several distinct regions. This map can be used to map materials onto the surface of objects and scenes (fig.˜4).
Creating 2D Scenes: The simple way to use the extracted map is to define every map region as a different material in an image and map different textures into different regions (Figure˜2.center.d). The resulting image is then overlayed on top of a random background image (Figure˜2.center.e). Shadow-like effects can be achieved by using another extracted soft map to darken some regions according to the map value. In regions where the value of the map is between 0 and 1, several textures are mixed according to the values of the map (fig.˜4).
Figure 4: Examples of 2D and 3D synthetic scenes (top) and their annotations (bottom). In the annotation, each material is marked as a different color. A mixture of materials marked as a mixture of their colors. The background is marked black. More samples can be seen at fig.˜11.
Creating 3D scenes: A more advanced approach is to use the generated maps to set PBR materials on object surfaces in 3D scenes using CGI software such as Blender[8]. This is done by creating 3D scenes and wrapping the extracted maps around objects’ surfaces in a process called UV mapping. Random PBR materials are then placed in each segment of the map (fig.˜2.left.d,e). Background, random objects, ground, and illumination are added to generate full 3D scenes with physics-based rendered materials (PBR), natural illumination, shadows, and curved surfaces (fig.˜1.center).
@@ 3.2 Extracting Textures and Materials
Materials in synthetic datasets are usually represented as SVBRDF/PBR, in which the spatial distribution of each material property (reflectivity, roughness, transparency, etc.) is given as a 2D map[6, 43]. Extracting textures and PBR materials from images involves first identifying regions with a uniform texture (which implies a uniform material, fig.˜2.right). This texture image is then used to extract maps corresponding to each material’s property.
@@ 3.2.1 Unsupervised texture extraction
Unsupervised texture extraction is traditionally done by finding image regions with a uniform distribution of color and gradients[57, 12, 42, 31, 62]. Finding these regions is done by splitting the image into square cells of 40x40 pixels each and identifying square regions where the distribution of the RGB values and their gradients is similar for all the cells in the region. Regions of more than 6x6 cells with similar distributions between all cells were extracted (fig.˜2.right). More accurately, a region was considered to have uniform texture if the similarity of the distributions between each pair of cells in the region had a Jensen-Shannon distance of less than 0.5. This similarity was tested separately for each of the RGB channels and their gradients. Finally, regions with too-uniform values or very high or low values are ignored, as they only contain smooth, textureless regions. It’s clear that this approach will fail to capture textures on bent surfaces or non-uniform illumination and shadows. However, the assumption is that for a large enough set of images, common material textures will appear in some cases in a simple setting and will be captured (section˜1.1). Once captured, they will be projected into scenes with different illuminations, and bent surfaces and will be learned for the general case. Some textures extracted using this approach are shown in figs.˜1 and 8.
@@ 3.2.2 Generating material property maps from texture images
Extracting PBR material maps from RGB textures requires guessing the physical properties of the material at each point of the texture image. There is no straightforward way to achieve this using an RGB image. Once again, the assumption will be that the properties of the material (reflectivity, roughness, and transparency) are correlated, in some cases, with simple image properties like color(section˜1.1), and for a common materials and a large enough set of images such correlation is bound to occur in some setting. Each material property map is generated by randomly choosing one of the six image channels (RGB or HSV), to represent this map (fig.˜2 left), image augmentation, thresholding, and random scaling/shifting were also used. Normal maps were generated by taking the gradient of the height/bumpiness map(fig.˜2.left). In some cases, the map was chosen to have a uniform random value. Some results are shown in figs.˜1 and 7. In addition, materials were also generated by mixing maps of existing PBR materials as described in appendix˜F.
@@ 4 Material State Segmentation Benchmark
Image collection and annotation: 1220 images for the benchmark were manually taken from a wide variety of real-world scenes (fig.˜3). Including food at different levels of cooking, liquids, solutes, foams, plants, rocks, minerals, sediments, grounds, landscapes, construction sites, metals, relics, fabrics, labs and processes like cooking, wetting, rotting, infections, corrosion, stains, and many others. The goal is to capture as much of the world as possible, both in terms of materials and environments. For materials with scatter, soft boundaries, and gradual transitions, it is often impossible to segment the entire region belonging to a material state. An alternative approach is to sample points in the image that represent the main materials’ states and their distribution. This allows annotators to focus on regions with clear annotation, but also to sample the harder and more complex regions without dealing with areas where the segmentation is unclear. The annotation procedure is as follows: 1) Pick points in each image that represent the distribution of each material state (fig.˜3). 2) Group the points according to their material state such that all points belonging to the exact same material state will have the same label or group (points of the same color in fig.˜3 are in the same group). 3) Assign relative similarity between groups. For example, if material B is a transition state between materials A and C, we can define the points in group B to be partially similar to both groups A and C, while groups A and C are not similar to each other. This circumvents the almost impossible task of assigning numerical similarity to material states (fig.˜3).
@@ 4.1 Evaluation metric
The evaluation approach is based on assessing the relative similarity between the points in the image. Every segmentation approach, soft or hard, can be easily converted to predict the similarity between two points in the image. For hard segmentation, this similarity will be one if the two points are in the same segment and zero if not. For evaluation, we use the triplet metric which goes as follows: a) Select three points in the image and define one as an anchor. b) Ask which of the remaining two points is more similar to the anchor according to the ground truth (GT) and according to the prediction; if the predictions and GT agree, we assign this as correct. c) If the two points have identical similarity to the anchor (according to the GT), we ignore this triplet. This is done for all sets of GT triplets in the image, and the average per material is used. For this metric, a 50% score means a random guess and a 100% score means perfect segmentation. The results are given in table˜1.
@@ 4.2 Additional benchmarks
While there are no other benchmarks for general zero-shot material state segmentation, there are a number of benchmarks that deal with subdomains of this task. These include segmentation of specific material states, like corrosion[48], minerals/ores[17, 16, 26, 28], dust[18], leaf disease[5], soil types and states[49, 50], microscopy[39, 40], or material phases (liquids, solids, foams, or powders)[24, 23]. Although each of these benchmarks is relatively narrow, when combined, they offer a way to cross-check the results of the main benchmark. Materialistic is a benchmark[53] that contains 50 images with different materials, not limited to specific classes. With materials segments that cover the whole object or regions with straight, smooth boundaries. It consists of distinct material types with no states or gradual transitions. Comparing the results of this benchmark to MatSeg offers a way to examine how much these elements affect segmentation accuracy. To evaluate the nets on these benchmarks, we select a random point in each material segment and use the nets to predict the similarity of each pixel in the image to the selected point. To turn this map into a hard segmentation map, we choose a threshold that maximizes the IOU between the predicted and GT masks. We use the mean IOU as the benchmark score. This is done with multiple points to obtain accurate statistics. The results are given in table˜2.
@@ 6 Results
Results for the MatSeg benchmark are given in table˜1 and fig.˜5. The net trained on the MatSeg data achieves good accuracy and significantly outperforms both SAM and Materialistic on both hard and soft cases. It can be seen from fig.˜5 that the net learned to predict highly scattered and complex patterns with soft and gradual transitions and achieves this for a wide range of domains. Hence, despite learning from synthetic data, the net manages to generalize for a wide range of real-world cases. In addition, it seems to be able to mitigate reflections and shadows fig.˜5. Both Segment Anything (SAM) and Materialistic nets performed far worse on the benchmark (table˜1). From fig.˜5, it seems that both SAM and Materialistic search for the boundary of the material based on the boundary of the object or bulk regions and struggle with scattered shapes and soft boundaries. SAM and Materialistic also seem to be less sensitive to the material state and more focused on matching materials based on coarse-grain type (fig.˜5). Both nets seem to search for bulk regions with smooth boundaries, which are likely to be more representative in their training set. SAM has failed to achieve high accuracy even after receiving 8 guiding input points (compared to 1 for other methods), suggesting that despite training on a vast amount of examples, it never encountered or learned such patterns. This seems to imply a major gap in the data collected for both SAM and Materialistic. This hypothesis is further supported by the fact that the SAM and Materialistic nets outperformed the MatSeg-trained net on the Materialistic test set (table˜2), which contains distinct materials that encompass either a full object or distinct regions with smooth hard boundaries. As such, their shape distribution resembles that of assets used in the Materialistic train set and the SAM polygon-based annotations. Hence, the MatSeg data seems to better model raw material states, while SAM and Materialistic seem to perform better on more structured scenes and object-guided segments.
Figure 5: Results: Predicted material similarity maps to a given point in the image. The input point is marked green in each panel. For SAM net the results are with 8 input points, 4 positive (marked green) and 4 negative (marked red). The materials similarity map is brighter for high similarity and dark for low similarity (similarity to the input point). More results can be seen at fig.˜10.
Table 1: Results on the MatSeg Benchmark (section˜4). The results are in triplet loss(section˜4.1). Points (pts) are the number of input points used to guide the segmentation. MatSeg 2D and 3D stand for nets trained on only 2D or 3D scenes, and Mixed stands for a net trained on a combination of 3D and 2D scenes. All stands for evaluation on all triplets. Soft stand for evaluation only on triplets with partial similarity between points(section˜4).
@@
Table 2: Results on various of datasets (section˜4.2). The results are in IOU for an optimal threshold (section˜4.2). Benchmarks with Scattered shapes and Gradual transitions are marked (S/G). Points (pts) is the number of input points used to guide the segmentation. The top one-shot (one point) method is marked.